### v1

# Optimal and Bounded Suboptimal Any-Angle Multi-agent Pathfinding

Abstract
Multi-agent pathfinding (MAPF) is the problem of finding a set of conflict-free paths for a set of agents. Typically, the agents’ moves are limited to a pre-defined graph of possible locations and allowed transitions between them, e.g. a 4-neighborhood grid. We explore how to solve MAPF problems when each agent can move between any pair of possible locations as long as traversing the line segment connecting them does not lead to the collision with the obstacles. This is known as any-angle pathfinding. We present the first optimal any-angle multi-agent pathfinding algorithm. Our planner is based on the Continuous Conflict-based Search (CCBS) algorithm and an optimal any-angle variant of the Safe Interval Path Planning (TO-AA-SIPP). The straightforward combination of those, however, scales poorly since any-angle path finding induces search trees with a very large branching factor. To mitigate this, we adapt two techniques from classical MAPF to the any-angle setting, namely Disjoint Splitting and Multi-Constraints. Experimental results on different combinations of these techniques show they enable solving over 30% more problems than the vanilla combination of CCBS and TO-AA-SIPP. In addition, we present a bounded-suboptimal variant of our algorithm, that enables trading runtime for solution cost in a controlled manner.
I INTRODUCTION
Multi-agent pathfinding (MAPF) is the problem of finding a set of conflict-free paths for a set of agents. Its applications include automated warehouses [1], traffic control [2], digital entertainment [3], etc. Different optimization variants of MAPF have proven to be NP-Hard [4], yet efficient optimal MAPF solvers have been proposed such as Conflict Based Search (CBS) [5], SAT-MDD [6], BCP [7] and others.
Most prior work focused on the classical version of the problem which makes many simplifying assumptions [8]. The most common of them are that i) the agents’ moves are limited to a given graph of possible locations and the transitions between them, and ii) all moves take one time unit. Indeed, these limitations are not well-suited for a variety of robotic applications and, thus, some works have begun to explore how to solve MAPF without them. In particular, the Continuous Conflict-based Search (CCBS) algorithm guarantees completeness and optimality while avoiding the need to discretize time and can handle agent actions with different durations [9]. Still, CCBS as well as many other previously proposed optimal MAPF solvers rely on the first assumption, i.e., that the agents move over a pre-defined graph of locations and allowed transitions between them.
We focus on relaxing this assumption, allowing each agent to move between any pair of locations – see Fig. 1. This type of pathfinding is often called any-angle pathfinding [10]. Algorithms such as Anya [11] and TO-AA-SIPP [12] have been proposed for optimal single-agent any-angle path finding, and a suboptimal any-angle MAPF solver was also proposed [13]. We propose AA-CCBS, the first any-angle MAPF algorithm that is guaranteed to return cost optimal solutions.
AA-CCBS integrates CCBS with TO-AA-SIPP. While it is guaranteed to return optimal solutions, AA-CCBS scales poorly since any-angle path finding induces search trees with a very large branching factor. To mitigate this, we propose several enhancements to AA-CCBS based on techniques from classical MAPF, namely disjoint splitting (DS) [14] and multi-constraints (MC) [15]. In particular, we suggest several novel variants of how MC can be applied for our any-angle setup, without compromising the theoretical guarantees.
We conduct a thorough empirical evaluation of different variants of AA-CCBS across standard benchmarks. The results show that the our enhancements AA-CCBS work very well, solving significantly more problems under a given time limit than vanilla AA-CCBS. We also show how AA-CCBS can be easily generalized to be a bounded-suboptimal algorithm, allowing a controlled way to trade off runtime for solution quality. Our experimental results demonstrate that allowing even a small amount of suboptimality allows solving many more problems.
Figure 1: Two optimal solutions of the same MAPF instance: the one composed of the cardinal moves only (left) and the one with any-angle moves (right). The cost of the latter is $22\%$ lower.
II RELATED WORKS
Lifting certain Classical MAPF [8] assumptions to make the resultant solvers more suitable to real-world applications has been studied recently. In [13] a prioritized any-angle MAPF planner was introduced and in [16] this planner was enhanced to support robots of different sizes and moving speeds. Other techniques for planning with kinematic constraints, including the post-processing of the plans produced by the classical MAPF solvers, were also proposed [17, 18, 19]. Yan and Li [20] proposed an involved three-stage solver based on Priority-Based Search [21], that handles not only kinematic constrainsts but accelerating and decelerating actions. All these solvers, unlike ours, are not optimal.
Optimal MAPF solvers that go beyond Classical MAPF setting also exist. A vast majority of them are adaptations of the seminal CBS algorithm [5]. In [22] individual roadmaps for the robots are constructed and then used by CBS (for collision detection each path in each roadmap is split into the uniform time resolution segments). A dedicated technique is suggested to decide when to re-construct these roadmaps. A combination of CBS with trajectory optimization methods was suggested to construct collision-free paths for a team of heterogeneous agents (quadrotors) [23, 24]. Kinodynamic variants of CBS were considered in [25, 26, 27]. In this work, we also leverage CBS, specifically its variant adapted to non-discretized timeline [9], to conduct multi-agent any-angle path finding with optimality guarantees.
III PROBLEM STATEMENT
Consider a set of vertices $V$ residing in the metric space, e.g., a grid where the vertices are the set of centers of the grid cells. There are $n$ designated start and goal vertices, and $n$ agents initially located at the start vertices. The agents are modeled as disks of a certain radius. Two types of actions for each agent are possible: wait at the current vertex (for an arbitrary amount of time) and move between the vertices. A move is valid only if the given line-of-sight function, $los:V\times V\rightarrow\{true,false\}$, returns $true$. This function checks whether an agent can follow a straight-line segment connecting the vertices without colliding with an obstacle, e.g., whether the set of grid cells swept by the disk-shaped agent contains only free cells. The cost of each action is its duration. The duration of a wait action may be arbitrary. The duration of a move action equals the distance between the vertices comprising this move (we assume that the agents start/stop instantaneously and move with the unit speed).
An individual plan for an agent is a sequence of timed actions $\pi=((a_{0},t_{0}),(a_{1},t_{1}),...,(a_{n},t_{n}))$, where $a_{i}$ is a move action, $t_{i}$ is the starting time of this action, and each action must start where the previous one ends. The wait actions are encoded implicitly in the plan, i.e. if $t_{i}+dur(a_{i})<t_{i+1}$ then the agent is waiting at the endpoint of $a_{i}$ for a time of $t_{i+1}-t_{i}-dur(a_{i})$. Here $dur(a_{i})$ denotes the duration of action $a_{i}$. The plans of distinct agents are said to be in conflict if there exists a time moment when the distance between the agents, executing these plans, is less than sum of their radii. The any-angle multi-agent pathfinding problem (AA-MAPF) asks to find a set of $n$ plans transferring the agents from their start vertices to the goal ones, such that each pair of plans is collision-free. The cost of the individual plan, $c(\pi)$, is the time when the agent reaches its goal. We wish to solve the problem optimally w.r.t. sum-of-cost objective, which is the sum of durations of the agents’ plans, $SOC=\sum_{i=1}^{n}{c(\pi_{i})}$.
IV BACKGROUND
Our planner is based on a combination of CCBS, [9] and TO-AA-SIPP [12]. Next, we briefly introduce these methods.
IV-A CCBS
CCBS is a variant of CBS [5] adapted to non-discretized timeline. It reasons over the continuous time-intervals instead of distinct isolated time steps. CCBS works by finding plans for each agent separately, detecting conflicts between these plans, and resolving them by replanning for the individual agents subject to specific constraints. A conflict between the two agents in CCBS is defined by $(a_{i},a_{j},t_{i},t_{j})$, representing that a collision occurs when executing the actions $a_{i}$ and $a_{j}$ at time moments $t_{i}$ and $t_{j}$, respectively. CCBS resolves a conflict by imposing a constraint on an agent and replan for that agent. The constraint imposed by CCBS is a tuple $(a,[t,t^{\prime}))$, stating that the agent is prohibited from taking action $a$ in the time range $[t,t^{\prime})$. The latter is called the unsafe interval. Several approaches were considered to compute $t^{\prime}$, the first time moment an agent can safely start executing the action, including closed-loop formulas [28].
To guarantee optimality, CCBS runs two search processes: a high-level search to choose which agent to constraint and a low-level search to find a path for that agent that satisfies its constraints. CCBS uses a best-first search for the high-level search and the Safe-Interval Path Planning (SIPP) algorithm [29] for the low-level search. SIPP is a specific adaptation of A* for space-time domains. It was originally proposed for single-agent pathfinding among dynamic obstacles, but can be easily adapted to handle CCBS constraints, as the latter can be viewed as dynamic obstacles that temporarily appear in the workspace.
IV-B TO-AA-SIPP
TO-AA-SIPP is a variant of SIPP that aims at finding an optimal any-angle path for an agent navigating in an environment with dynamic obstacles. As in regular SIPP, its search node is identified by a tuple $(v,[t_{l},t_{u}])$, where $v$ is a graph vertex and $[t_{l},t_{u}]$ is a safe time interval that dictates when the agent may safely reside at $v$.
The principal difference between TO-AA-SIPP and SIPP is that the former does not iteratively build a search tree by expanding search nodes. Instead, it generates all search nodes beforehand (this is possible due to the usage of time intervals as the nodes’ identifiers) and iteratively tries to identify the correct parent relationships between the nodes to obtain the optimal solution. The authors refer to this technique as “inverted expansions”. TO-AA-SIPP was proven to be complete and optimal.
V AA-CCBS
To solve AA-MAPF, we propose to use TO-AA-SIPP as the low-level planner for CCBS. We call this algorithm AA-CCBS. Unfortunately, AA-CCBS struggles to solve even simple instances. E.g., the small AA-MAPF instance with 3 agents depicted in Fig. 1 requires expanding $6,660$ high-level nodes and consequently a similar number of low-level calls to TO-AA-SIPP. To enable AA-CCBS to scale more gracefully, we propose two vital enhancements: Multi-Constrains (MC) and Disjoint Splitting (DS).
V-A Multi-Constraints in Any-Angle MAPF
CCBS resolves a conflict by adding a single constraint to one of the agents and replanning. However, adding multiple constraints (multi-constraint) when resolving a single conflict can reduce the number of high-level search iterations [30, 15]. In particular, Walker et al. [15] proposed the Time-Annotated Biclique (TAB) method for adding multiple constraints in domains with non-uniform actions.
For two conflicting actions, $(a_{i},t_{i})$ and $(a_{j},t_{j})$, TAB iterates over all actions that have the same source vertices as either $a_{i}$ or $a_{j}$ and identifies the subsets of them, $A_{i}$ and $A_{j}$, that are mutually conflicting, i.e. each pair of actions from $A_{i}\times A_{j}$ lead to a conflict (if they start at $t_{i}$, $t_{j}$ respectively). The multi-constraint (MC) added to the first agent comprises the constraints $(a_{i},[t,t^{\prime}])$ where $[t,t^{\prime}]$ is the largest time interval that is fully included into the unsafe intervals induced by $a_{i}$ and all $a_{j}\in A_{j}$. The MC added to the second agent is defined similarly. TAB is applicable in AA-CCBS. We call this variant MC1.
Figure 2: An example of vanilla AA-CCBS multi-constraint.
Example. Consider two actions sets $\{a_{i_{1}},a_{i_{2}}\}$, $\{a_{j_{1}},a_{j_{2}}\}$ and time moments $t_{i},t_{j}$. Let $[t_{i},t^{\prime}]=unsafe(a_{i_{1}},a_{j_{1}})\cap unsafe(a_{i_{1}},a_{j_{2}})$ and $[t_{i},t^{\prime\prime}]=unsafe(a_{i_{2}},a_{j_{1}})\cap unsafe(a_{i_{2}},a_{j_{2}}])$, where $unsafe$ denotes the unsafe interval of $a_{i_{k}}$ starting at $t_{i}$ w.r.t. $a_{j_{l}}$ starting at $t_{j}$. Then, the MC associated with the first action set is $\{(a_{i_{1}},[t_{i},t^{\prime}));(a_{i_{2}},[t_{i},t^{\prime\prime}))\}$. It dictates that the agent is not allowed to perform either $a_{i_{1}}$ or $a_{i_{2}}$ in the respective time intervals. Fig. 2 shows an illustrative example.
The problem with MC1 is that in AA-MAPF the number of mutually-conflicting actions may be large. Thus, after intersecting their unsafe intervals the final unsafe interval for a certain action is likely to get trimmed. For example, consider action $a_{1_{1}}$ of Fig. 2, which has an unsafe interval of $[0;0.522)$ in the MC, despite its unsafe interval w.r.t $a_{2_{1}}$, ending at $1.106$, is twice longer. A shorter unsafe interval means that the power of constraint diminishes. To mitigate this issue we suggest two modifications.
First, we suggest to consider only a limited set of actions to form $A_{i}$ ($A_{j}$ similarly). The source of each action should coincide with the source of the initially conflicting action $a_{i}$. The destination must be a vertex that is swept by the agent when executing $a_{i}$, i.e. the agent’s body intersects this vertex (grid cell). Thus the resulting actions form a “stripe” along $a_{i}$ – see Fig. 3 (the cells that form a stripe are highlighted).
The second modification is that we filter out the actions from the MC that lead to trimming the original unsafe interval. In particular, if $unsafe(a_{i},a^{\prime}_{j})\not\subset unsafe(a_{i},a_{j})$ then $a^{\prime}_{j}$ is excluded from $A_{j}$. Here $a_{i}$ and $a_{j}$ are the original conflicting actions, and $\{a^{\prime}_{j}\}$ is the set of actions that are in conflict with all $a\in A_{i}$. The combination of those two enhancements is dubbed MC2.
The detailed pseudo-code of constructing MC2 is shown in Alg. 1. After initialization (Lines 1-3), we identify the swept cells (Line 4) and form the candidate sets of actions $\hat{A_{i}},\hat{A_{j}}$ (Line 5). Then, in Line 6, we merge them into $\hat{A}$ in such a way that the actions from $\hat{A_{i}}$ and $\hat{a_{j}}$ alternate (this merging procedure is denoted as $zip$ in code). Next, we iterate over $\hat{A}$ and filter out actions that lead to trimming the unsafe interval of $a_{i}$ (Lines 10-12). After pruning the candidate sets we iterate over the actions once again to compute their unsafe intervals (Lines 15-21). The resulting sets of action-interval pairs form MC2.
Figure 3: Actions comprising different versions of multi-constraints. Please note that the time intervals of the actions in MC2 and MC3 are notably larger compared to MC1. Thus MC2 and MC3 are expected to exhibit more pruning power.
Input: Conflicting actions (and times) $a_{i},t_{i},a_{j},t_{j}$
Output: A pair of multi-constraints $\{A_{i},A_{j}\}$
1
2$[t_{i},t^{\prime})\leftarrow$ unsafe interval of $a_{i}$ w.r.t. $(a_{j},t_{j})$;
3 $[t_{j},t^{\prime\prime})\leftarrow$ unsafe interval of $a_{j}$ w.r.t. $(a_{i},t_{i})$;
4 $A_{i}\leftarrow\emptyset$; $A_{j}\leftarrow\emptyset$;
5 $C_{\{i,j\}}\leftarrow$ vertices swept by $a_{\{i,j\}}$;
6 $\hat{A}_{\{i,j\}}\leftarrow$ valid actions that start at the source of $a_{\{i,j\}}$ and end in $C_{\{i,j\}}$;
7 $\hat{A}\leftarrow zip(\hat{A_{i}},\hat{A_{j}})$;
8
9for each $\hat{a}\in\hat{A}$ do
10       if $source(\hat{a})=source(a_{i})$ then
11             if IsConflict($\hat{a},A_{j}$)=true then
12                   $[t_{j},\hat{t})\leftarrow$ unsafe interval of $a_{j}$ w.r.t. $(\hat{a},t_{i})$;
13                   if $t^{\prime\prime}>\hat{t}$ then
14                         remove $\hat{a}$ from $\hat{A_{i}}$;
15
16
17
18      else
19             handling $a_{j}$ similarly (Lines 9-12);
20
21
22
23for each $a\in\hat{A_{i}}$ do
24       $t_{min}\leftarrow\infty$;
25       for each $a^{\prime}\in\hat{A_{j}}$ do
26             $[t_{i},t_{cur})\leftarrow$ unsafe interval of $a$ w.r.t. $(a^{\prime},t_{j})$;
27             $t_{min}\leftarrow\min\{t_{min},t_{cur}\}$;
28
29      $A_{i}\leftarrow A_{i}\cup\{(a,[t_{i},t_{min}))\}$;
30
31
32Process $\hat{A_{j}}$ similarly (Lines 15-20);
33
34return $\{A_{i},A_{j}\}$
Algorithm 1 Forming MC2 for AA-CCBS
Next, we suggest one more way of composing the action sets of MC, when we include into the constrained set actions that start at the different graph vertices. This is possible as by definition the set of actions comprising $A_{i}$ and $A_{j}$ must be mutually conflicting but the definition does not specify what source and target vertices of these actions should be. In specifics, we enlarge the action sets of MC2 with the actions that have the same target vertex (as the original conflicting actions) and the source vertices that lie on the same stripe as before – see Fig. 3 (right). The rationale is that these actions are likely to lead to collisions that will happen between the same agents and nearly in the same place. Thus it is natural to include them in the same multi-constraint. The pseudo-code for constructing MC3 largely repeats the one for MC2 (with certain additions) and is omitted for the sake of space.
Figure 4: The results of evaluation of all the versions of AA-CCBS. a) shows the amount of totally solved instances b) show the time required to find a solution c) demonstrates the comparison of the three best-performing methods in terms of HL expansions.
V-B Disjoint Splitting
Disjoint splitting (DS) is a powerful technique for reducing search effort in CBS [14] that was shown to be effective in CCBS as well [31]. In DS for CCBS, a conflict $(a_{i},a_{j},t_{i},t_{j})$ is resolved in two ways: one CCBS child gets a regular (negative) constraint on agent $i$, while the other one - a positive constraint to agent $i$ and a negative constraint to agent $j$. The positive constraint comes in the conventional CCBS form of $(a,[t,t^{\prime}))$ but dictates the agent must perform action $a$ in a time moment belonging to $[t,t^{\prime})$. Thus action $a$ becomes a landmark in the sense that the low-level search has to, first, find a plan to the source of $a$, then execute $a$, and then plan to the goal (or to the next landmark).
In this work, we suggest two options for implementing DS for AA-CCBS: using DS as-is and using DS with MC. The first one is straightforward, while the second requires more attention.
To integrate DS with MCs that start at the same vertex but end in different ones, like MC1 or MC2, we need to impose a combined positive constraint that will dictate an agent to perform any of the actions comprising the constrained set. Thus, we need to reason over the multiple distinct start locations for the consecutive search, which is not trivial. Integrating DS with MC3 is even more involved as here we need to handle landmarks that are composed of the actions that start at different locations.
To simplify the integration of DS and MC in AA-CCBS we suggest to avoid creating landmarks out of multi-constrains, while still operating with conventional negative multi-constraints. In particular, consider a conflicting pair of timed actions, $(a_{i},t_{i},a_{j},t_{j})$. As in the original DS, we impose a singular negative constraint on the first agent in the left CCBS child and the corresponding singular positive constraint in the right child. Additionally, the second agent gets a negative multi-constraint in the right child, which is computed with respect to the action $(a_{i},t_{i})$ only. In this way, there is no need to handle positive MCs, while the regular (singular) positive constraint is enforced with the negative MC (in one of the descendants in CCBS search tree).
Theorem 1
AA-CCBS with any of the enhancements — MC1, MC2, MC3, and DS — is sound, solution-complete, and optimal, i.e., it is guaranteed to return a valid solution is such exists, and the solution it returns is cost-optimal.
Proof outline: Soundness, solution-completeness, and optimality for vanilla AA-CCBS is trivial, following the proofs of CCBS [9] and TO-AA-SIPP [12]. To prove that the proposed improvements (MC1, MC2, MC3, and DS) preserve these properties, it is sufficient to show that AA-CCBS with them always resolves conflicts with a sound pair of constraints [32]. A pair of constraints is sound if in any valid solution, at least one of these constraints is satisfied. That is, any solution in which both constraints are violated is not valid. In MC1, MC2, and MC3, by construction, for every pair of multi-constraints $A_{i}$ and $A_{j}$ they return it holds that any pair of constraints $(a_{i}^{*},T_{i}^{*})\in A_{i}$ and $(a_{j}^{*},T_{j}^{*})\in A_{j}$ form a sound pair of constraints. Thus, any solution that violates a constraint in $A_{i}$ and a constraint in $A_{j}$ cannot be valid. A similar argument holds for DS due to the negative constraints it imposes.
VI Experimental Results
We implemented AA-CCBS in C++ [^1] https://github.com/PathPlanning/Continuous-CBS/tree/AA-CCBS and conducted the experimental evaluation on five different maps from the well-known in the community MovingAI benchmark [8]: empty-16-16, random-32-32-20, maze-32-32-4, den321d and warehouse-10-20-10-2-2. For each map, the benchmark provides 25 scenarios. Each scenario is a set of start-goal locations. We take two agents, run the algorithm, and then increment the number of agents until the algorithm is not able to find a solution within a time-limit of 300 seconds. In the latter case, we proceed to the next scenario. We evaluate six versions of AA-CCBS: AA-CCBS (vanilla), AA-CCBS+MC1, AA-CCBS+MC2, AA-CCBS+MC3, AA-CCBS+DS, AA-CCBS+DS+MC3.
Fig. 4 shows the results. The left pane shows the number of solved instances (with the breakdowns for each map). The results indicate that MC1 and MC2 do not provide significant improvement over vanilla AA-CCBS. MC3, DS, and DS+MC3, however, notably outperform it, especially the ones that use DS. The central pane shows the number of tasks solved within a certain time cap. The X-axis shows the number of tasks and Y-axis shows the time needed to solve this amount of tasks. One can note that MC3 outperforms all other variants of AA-CCBS for easy tasks, i.e. for the ones that need less than 0.1 seconds to be solved. Meanwhile DS variants of AA-CCBS clearly outperform MC3 when the time, needed to solve an instance, is greater than 1 second. The right pane of Fig. 4 depicts the number of the high-level iterations of AA-CCBS with MC3, DS, DS+MC3 for each MAPF instance that was solved by all these methods. As seen on the plot, this number is lower for MC3 for nearly half of the instances. This also confirms that MC3 solves some (i.e. easy) instances faster than DS and MC2+DS, while it is outperformed on the other (hard) ones.
Overall the obtained results confirm that both MC and DS are valuable techniques to enhance AA-CCBS performance. MCs (specifically, MC3) are well-suited for the easy instances while the DS – for the harder ones. Generally, we would recommend to use AA-CCBS+DS+MC3 to solve any-angle MAPF optimally.
Figure 5: The results of the evaluation of AA-CCBS and its modifications with different suboptimality factors.
VII Bounded-Suboptimal AA-CCBS
Given a real number $w>1$ a bounded suboptimal (BS) solution is the one whose cost does not exceed the cost of the optimal one by more than a factor of $w$. Finding a BS solution is often much easier and faster than finding an optimal one. Creating a BS variant of AA-CCBS with any of our enhancements is straightforward: replacing the best-first search in the AA-CCBS high-level with a Focal search [33]. [^2] Creating a BS low-level search is not trivial in our case and goes beyond the scope of this paper. I.e. at each high-level iteration we, first, form a subset of nodes whose cost does not exceed the cost of the most promising node (the one which regular AA-CCBS would have picked) by a factor of $w$ and, second, choose from this subset the node with the smallest number of conflicts breaking ties in favor of nodes containing more constraints. Thus we force the search to more greedily resolve the conflicts, while still maintaining the bound on the suboptimality of the resultant solution.
We implemented such a BS version of AA-CCBS with the three best-performing enhancements, DS, MC3, DS+MC3, and evaluated them on the same problem instances as before. Fig. 5 shows the results. In the upper part the box-and-whiskers plot of the resultant solution cost depending on the suboptimality bound are shown. The table at the bottom shows the number of the solved instances across all maps.
Indeed, MC3 notably increases the performance of AA-CCBS when searching for bounded suboptimal solutions (i.e. when $w>1$), while DS variants are not even able to outperform vanilla AA-CCBS. We hypothesize that this is due to the positive constraints imposed by DS do not lead to immediate eliminating the conflict that causes these constraints. When searching for optimal solutions this is compensated by reducing the branching factor. However the search for suboptimal solutions is more greedy by design and branching factor becomes much less important. Thus, DS looses much of its power, while retaining the drawback of not immediately eliminating the conflict. At the same time, MC3 imposes constraints on a larger number of actions, and as the search greedily goes deeper it actually eliminates the conflict. Moreover, due to the requirement to satisfy all the imposed positive constraints, the actual costs of the solutions found by DS versions are on average more than twice higher than the ones of MC3. The actual overhead in solution costs of MC3 version is not higher than 2% on average even when the suboptimality bound is $w=1.25$. The obtained results provide further support that MC3 is a useful technique: not only does it speed the search for optimal solutions on easy instances, but it also significantly increases the performance of bounded-suboptimal AA-CCBS.
VIII CONCLUSION AND FUTURE WORK
In this work, we presented AA-CCBS, the first optimal any-angle MAPF algorithm. We showed how to incorporate existing CCBS enhancements, namely Disjoint Splitting and Multi-Constraints, into AA-CCBS to allow it to scale better. Specifically, we introduced three ways to implement Multi-Constraints, providing stronger pruning of the high-level search tree. AA-CCBS can also be easily adapted to return bounded-suboptimal solutions. Experimental results confirm that the suggested enhancements have a huge impact on the algorithm’s performance. Future work can explore more sophisticated procedures of forming multi-constraints, identifying which sets of actions should be considered in each step, as well as adapting incremental search techniques in the any-angle low-level search.